OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today?

OpenStudy (anonymous):

You've got a system of equations \[x+y=1070\]\[3x-2y=110\] where x=#of one step equations where y=# of 2 step equations go from there

OpenStudy (anonymous):

@doulikepiecauseidont im still confused

OpenStudy (anonymous):

Ok well we can call the number of 1 step equations x and the number of 2 step equations y, since we don't what they are (thats why we use variables) anyways it said that the number of them combined is 1070 so thats how we got the first equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then they said 3 times x (the number of 1 step equations - 2 times y(the number of 2 step equations)=110 which is where we get the second equation from

OpenStudy (anonymous):

Now the question is askking for the number of 1 step equations and we represented that by x, so basically we're solving for x, understand so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i got 450

OpenStudy (anonymous):

Ok well theres a way to solve for x and y if they're equal to the same thing and we have two equations for them, it's called solving for a system of equations

OpenStudy (anonymous):

Solving for the 1st equation for x: \[x+y=1070\]\[x=1070-y\]

OpenStudy (anonymous):

x=450 y=620

OpenStudy (anonymous):

Oh, ok you got it! I'm just used to having to explain how to solve for it lol.

OpenStudy (anonymous):

haha thanks

OpenStudy (anonymous):

so just to be shure the answer is 450

OpenStudy (anonymous):

Whay kind of math are you doing anyways? "Math Idol judges"? smfh and yeah thats it, you can check by plugging the numbers in, which is what I did with your answers (shortcut to actually solving it for me) :P

OpenStudy (anonymous):

algebra 1